• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Puckett

    Log entry time 08:01:00 on April17,2007

    Entry number 199351

    keyword=Owl Shift Summary

    Shift Workers:
    A. Puckett
    E. Khrosinkova

    Running 100 muA on 3He.

    Run summary(run #, start-stop, N, replay?)
    2555, 00:01-00:27, 1.2M, Y
    2556, 00:29-00:57, 1.2M, Y
    2557, 00:58-01:22, 1.2M, Y
    2558, 01:23-01:36, 465K, Y

    run 2558 ended early because we lost beam. MCC is going to do their QE measurement, and hopefully beam is back soon.

    At ~1:47 the beam is back and we are rolling again with run 2559. Elena changed ADC thresholds, see her HALOG entry.

    2559, 01:49-02:16, 1.2M, Y
    2560, 02:18-02:42, 1.2M, Y
    2561, 02:43-03:11, 1.2M, Y
    2562, 03:12-03:39, 1.2M, Y
    2563, 03:40-04:06, 1.2M, Y
    2564, 04:07-04:21, 600K, Y

    Beam has been off for a few minutes, we ended the run, and are waiting for it come back up. Will call MCC in a few more minutes.

    MCC calls to notify us that there was a problem with an RF cavity and beam was down for ~25 minutes to fix the problem.

    2565, 04:43-05:10, 1.2M, Y
    2566, 05:10-05:36, 1.2M, Y
    2567, 05:37-06:06, 1.2M, Y
    2568, 06:07-06:36, 1.2M, Y
    2569, 06:37-07:03, 1.2M, Y
    2570, 07:04-07:31, 1.2M, Y
    2571, 07:32-07:57, 1.2M, Y
    2572, 07:58-??

    End of shift. Smooth shift. Where is the day shift?